What is the centripetal acceleration of an object going in uniform circular Motion with a radius of 3m and a velocity of 5m/s


1
Expert's answer
2021-02-12T16:16:57-0500
"a_c=\dfrac{v^2}{r}=\dfrac{(5\ \dfrac{m}{s})^2}{3\ m}=8.33\ \dfrac{m}{s^2}."